OpenStudy (anonymous):

Which function has the greatest rate of change? A. 6x – 3y = 12 B. y = 5x + 4 C. x y 1 2 3 8 5 14

OpenStudy (anonymous):

@iGreen can you help?

OpenStudy (igreen):

Rate of change is slope.. y = 5x + 4 This is in the form of y = mx + b, where 'm' is slope..so what's the slope?

OpenStudy (anonymous):

5

OpenStudy (igreen):

Yep, now for the first one, we have to change it to the form y = mx + b. 6x - 3y = 12 Subtract 6x to both sides: -3y = -6x + 12 Divide -3 to both sides: y = 2x - 4 What's the slope here?

OpenStudy (anonymous):

2

OpenStudy (igreen):

Yes, now for the 3rd one we can take any of the two points and plug them in the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{8-2}{3-1}\) Simplify that to find the slope

OpenStudy (anonymous):

\[\frac{ 6 }{ 2 } = 3\]

OpenStudy (igreen):

Yep, you got it. So which one has the greatest slope?

OpenStudy (anonymous):

B?

OpenStudy (igreen):

A had a slope of 5 B had a slope of 2 C had a slope of 3

OpenStudy (anonymous):

what about D?

OpenStudy (igreen):

It has a slope of 4.

OpenStudy (anonymous):

ok so its A

OpenStudy (igreen):

Yes,